The sum of Greg's age and Sherri's age is 42. Greg's age 14 years from now will be twice what Sherri's age was 11 years ago. What are their ages now?
Thank you!

Answer by lwsshak3(11628)

The sum of Greg's age and Sherri's age is 42. Greg's age 14 years from now will be twice what Sherri's age was 11 years ago. What are their ages now?
***
let x=Greg's age
42-x=Sheri's age
..
14 yrs from now:
x+14=Greg's age
..
11 yrs ago:
42-x-11=31-x=Sheri's age
..
x+14=2(31-x)
x+14=62-2x
3x=48
x=16
42-16=26
Greg's age=16
Sheri's age=26